OpenStudy (anonymous):

emily purchased 6 tokens for an amusement park ride. Each token is worth $0.75. She wants to ride the roller coaster which costs $3.00 per ride. What fractional part of her tokens will she have to use for the ride? Just for fun;)

jimthompson5910 (jim_thompson5910):

It costs $3 a ride and each token is worth $0.75 So she needs 3/0.75 = 4 tokens Therefore, the fractional part is 4/6 = 2/3